OpenStudy (geometry_hater):

For what value of k does the equation kt^2 +5t = 0 have exactly one real solution for t?

jimthompson5910 (jim_thompson5910):

Hint: ax^2+bx + c = 0 has exactly one solution for x when b^2 - 4ac = 0

OpenStudy (anonymous):

kt^2 +5t = t(kt +5) = 0 already has t = 0 as a solution, regardless of k

OpenStudy (noelgreco):

...but since c=0 and b=5 no can do. k=0 works